# LlavaGuard: An Open VLM-based Framework for Safeguarding Vision Datasets and Models

### v2

@@ Appendix I LlavaGuard Dataset
The dataset consists of 5,466 unique samples (3,242 safe and 2,224 unsafe). Out of these, 3,242 samples are based on the default policy, while the remainder use augmented policies. The rationales for each sample are generated using Llava-34B. App. Fig. 10(a) provides an overview of the dataset, along with additional insights into the distribution of safety categories and ratings. The dataset is split into 4571 for training, 71 for evaluation, and 824 for test. The test set is balanced across safety categories and ratings (cf. App. Fig. 10(b)). To achieve an equal distribution of safe and unsafe samples during training, we apply oversampling to the unsafe samples in the training set, ultimately leading to a total of 5592 samples. Importantly, no images from the test set are observed during training. We make our annotated dataset and pipeline publicly available to stimulate further research.
@@
For the user study, we involved colleagues who are not safety domain experts but can deal with safety-related content. The study included three experts, each annotating 105 images along with LlavaGuard assessments. This group brought diverse backgrounds to the evaluation. The images are randomly selected from a mix of real and synthetic datasets: CC12M (Changpinyo et al., 2021b), COCO (Lin et al., 2014b), ImageNet (Deng et al., 2009), GenAI-Bench (Li et al., 2024a), and Stylebreeder (Zheng et al., 2024).

### v3

@@ Appendix I LlavaGuard Dataset
The dataset is annotated by the authors. All annotators are male, White, and between 20-40 years old. We adopted a prescriptive annotation approach, collaboratively developing a taxonomy that defines a detailed categorization. For edge cases, all annotators discussed potential contradictions to achieve consensus.
The dataset consists of 5,466 unique samples (3,242 safe and 2,224 unsafe). 3,242 samples of the dataset are based on the default policy, while the remainder use augmented policies. The rationales for each sample are generated using Llava-34B. App. Fig. 10(a) provides an overview of the dataset, along with additional insights into the distribution of safety categories and ratings. The dataset is split into 4571 for training, 71 for evaluation, and 824 for test. The test set is balanced across safety categories and ratings (cf. App. Fig. 10(b)). To achieve an equal distribution of safe and unsafe samples during training, we apply oversampling to the unsafe samples in the training set, ultimately leading to a total of 5592 samples. Importantly, no images from the test set are observed during training. We make our annotated dataset and pipeline publicly available to stimulate further research.
@@
For the user study, we involved colleagues who are not safety domain experts but can deal with safety-related content. For the user study, participants were aged 20-30, 30% male and 70% female, with 70% identifying as White (from the US and Europe) and 30% as Asian-American. The study included three experts, each annotating 105 images along with LlavaGuard assessments. This group brought diverse backgrounds to the evaluation. The images are randomly selected from a mix of real and synthetic datasets: CC12M (Changpinyo et al., 2021b), COCO (Lin et al., 2014b), ImageNet (Deng et al., 2009), GenAI-Bench (Li et al., 2024a), and Stylebreeder (Zheng et al., 2024). Despite the demographic differences between dataset annotators and user study participants, we observed a high level of agreement.
@@
Figure 13: False-positive examples from the LlavaGuard test set. Shown are images that LlavaGuard-7B has incorrectly classified as unsafe, with common failure modes including decision-boundary ambiguity and misinterpretation of embedded text. For example, signage prohibiting inappropriate behaviors was erroneously flagged as violating safety policies.
Appendix K Failure Case Analysis
Upon analysis of false positives, we identified several recurrent failure modes in LlavaGuard. Images situated near decision boundaries were particularly challenging for the model to classify. Furthermore, LlavaGuard occasionally struggled to interpret embedded textual elements within images. For instance, images containing signage explicitly prohibiting sexual harassment of nurses were misclassified as unsafe content. Illustrative examples of such errors are presented in Fig. 13.